### v2

# What Media Frames Reveal About Stance: A Dataset and Study about Memes in Climate Change Discourse

Abstract
Media framing refers to the emphasis on specific aspects of perceived reality to shape how an issue is defined and understood. Its primary purpose is to shape public perceptions often in alignment with the authors’ opinions and stances. However, the interaction between stance and media frame remains largely unexplored. In this work, we apply an interdisciplinary approach to conceptualize and computationally explore this interaction with internet memes on climate change. We curate ClimateMemes, the first dataset of climate-change memes annotated with both stance and media frames, inspired by research in communication science. ClimateMemes includes 1,184 memes sourced from 47 subreddits, enabling analysis of frame prominence over time and communities, and sheds light on the framing preferences of different stance holders. We propose two meme understanding tasks: stance detection and media frame detection. We evaluate LLaVA-NeXT and Molmo in various setups, and report the corresponding results on their LLM backbone. Human captions consistently enhance performance. Synthetic captions and human-corrected OCR also help occasionally. Our findings highlight that VLMs perform well on stance, but struggle on frames, where LLMs outperform VLMs. Finally, we analyze VLMs’ limitations in handling nuanced frames and stance expressions on climate change internet memes.
What Media Frames Reveal About Stance:
A Dataset and Study about Memes in Climate Change Discourse
[…]
1 Introduction
(a) convinced stance with and impact frames
(b) skeptical stance with hoax frame
Figure 1: Two climate change memes conveying opposite stances using different media frames.
Memes are a powerful communication format in online discourse that reflect communities’ cultural and social dynamics  Davis et al. (2016); Zhang and Pinto (2021). Memes are multimodal digital items that combine images and texts to convey complex viewpoints in a compact and engaging format Sharma et al. (2020); Liu et al. (2022). Memes express the communicator’s position towards a particular topic, i.e., stances as Mohammad et al. (2016) defines, through the specific combination of modalities such as text and images.
While stance reflects the creator’s opinion toward a target, the specific narrative used to convey a certain stance is shaped by media frames. Media framing refers to selecting specific aspects of a perceived reality in communication to portray how an issue is defined, how its causes are interpreted, how its moral implications are evaluated, and what potential solutions are considered Entman (1993); Gidin (1980). Depending on their stance, creators may gravitate toward different framing strategies Snow and Benford (1992). However, the interaction between stance and media frames remains under-studied, particularly in their representation through humorous social media content, such as memes, and its impact on debates of global significance, such as the climate change.
Climate change (CC) memes are a vital component of social media, such as Twitter/X Ross and Rivers (2019). For example, Figure 1(a) conveys a convinced stance towards CC by using and impact frames (further detailed in §3.2) to affirm the evidence of global warming and its disheartening consequences. Conversely, Figure 1(b) conveys a skeptical stance using the hoax frame, claiming that CC is not a major issue or even not real and suggests that politics may distort the CC issue.
In this paper, we analyze stances and media frames in CC memes111 We will release the data and code with the camera-ready submission. by examining the following three research questions (RQs):
  • RQ1: How do different media frames shape the visual representation of climate change in memes across varying stances? We introduce ClimateMemes, a dataset of CC memes, consisting of 1,184 CC memes from 47 subreddits, manually annotated with stance on climate change and the media frames they invoke (§3) to analyze how memes convey stance through strategic media framing (§4).
  • RQ2: Can state-of-the-art VLMs and LLMs accurately detect stances conveyed by memes and the corresponding media frames? We extend stance detection from text
    and propose a new task of multi-label media frame detection on CC memes. We evaluate two open-source VLMs and their backbone LLMs (§5.1) and investigate the effects of few-shot experiments and input modalities on these two tasks (§5.2). We found that while synthetic meme captions cannot yet fully replace human-annotated ones, they still improve the VLMs’ performance on both tasks. Yet, LLMs outperform VLMs on frame detection.
  • RQ3: Can taxonomies from communication science provide more insights on stance and media frame detection results? We recruit communication science specialists to annotate humor type, person, and responsibility features on 235 test CC memes. Our analyses reveal that the performances of VLMs and LLMs degrade markedly on memes that are jokes, about political figures, and individual (micro-level) responsibilities (§6).
2 Background
Figure 2: ClimateMemes’s pipeline of data collection, filtering, and annotations of stance, media frames, etc.
2.1 Memes
Internet memes are multimodal and humorous forms of expression that are popular across various digital channels, especially on social media Shifman (2014). They often use replicated and modified templates and are circulated among users to convey new, context-specific meanings. For example, Figure 1(a) uses the “Simpsons so far” template to highlight the ongoing trend of global warming. In controversial political arenas, such as the climate discourse, internet memes are seen as an effective tool for capturing attention, allowing users to communicate their stances through impactful imagery and humor Ross and Rivers (2019).
Nguyen and Ng (2024) categorize meme understanding into three main types: classification, interpretation, and explanation. Classification aims to assign labels to memes, such as identifying harmful content  Kiela et al. (2020); Pramanick et al. (2021); Cao et al. (2022); Hee et al. (2023), sentiment  Sharma et al. (2020); Chauhan et al. (2020), or figurative language Liu et al. (2022); Xu et al. (2022). Interpretation tasks focus on understanding and generating insights from memes, such as generating captions or analyzing the metaphor between the image and text componants Hwang and Shwartz (2023); Chen et al. (2024). Explanation tasks go a step further by generating textual justifications for the labels assigned to memes (Hee et al., 2023). In this study, we curate ClimateMemes and introduce two meme understanding tasks: stance detection and media frame detection. We also collect human-corrected OCR and meme captioning as a basis for future tasks.
2.2 Media Frames
Strategic media framing refers to the selective presentation of information to influence audience attitudes or evoke specific reactions Snow and Benford (1992). Social and communication science research has relied on framing concepts for analyzing how information is selected and presented in the media. Scholars in the field have created codebooks for manual identification of generic and issue-specific frames in media contexts.
The Media Frames Corpus (Card et al., 2015), focusing on three specific issues: immigration, smoking, and same-sex marriage, brought the methodologies of framing into our NLP community. Subsequent efforts have expanded this foundation, including proposals for general, issue-independent frame taxonomies (Johnson et al., 2017), computational framing analysis approaches (Mendelsohn et al., 2021; Ali and Hassan, 2022), and highlighting the importance of cognitive, linguistic, and communicative aspects beyond topical content in frame detection (Otmakhova et al., 2024).
In the context of climate change, framing has been studied to understand its role in public discourse and media representation. Stede et al. (2023) utilize generic frames, which are more abstract and commonly observed across political discussions, to analyze climate change in Nature and Science editorials. Chen et al. (2022) study how frames evolve within public events, emphasizing their divergence and convergence in shaping climate change narratives, while Frermann et al. (2023) analyze how news articles across the political spectrum frame climate change. To the best of our knowledge, this paper presents the first dataset of multimodal memes annotated with media frames and analyzes how frames interact with stances.
3 ClimateMemes Dataset
This section describes ClimateMemes, a dataset of 1,184 CC memes from 47 subreddits annotated with media frames and stances. Figure 2 illustrates our data processing pipeline. We discuss meme collection and climate filtering (§3.1), and present guidelines for stance and frame annotations (§3.2).
3.1 Source Memes and Climate Filter
Data Source.
To collect CC memes, we search subreddits with “meme” in their names and filter the topic of posts with the keyword “climate”. To obtain diverse climate change perspectives, our collection includes subreddits like r/ClimateMemes (primarily hosting climate activists) and r/ConservativeMemes (reflecting a more skeptical community on CC).
Out of 2,015 initially collected images, 1,184 CC-associated memes from 47 subreddits remained after filtering. Table 1 shows the top 6 subreddits that contribute to 79.6% of CC-associated memes (see Appendix A for a complete list of subreddits). The table also presents distributions of stance and frame labels (to be detailed in §3.2).
r/subreddit #m conv./skep./nei. #f top 3 frames
ClimateMemes 591 94.1 /  3.2 / 2.7 2.35 adeq, caus, impa
TheRightCantMeme 90 13.2 / 83.5 / 3.3 1.70 hoax, prop, caus
dankmemes 90 82.3 / 13.3 / 4.4 1.84 adeq, impa, real
memes 76 92.1 /  1.3 / 6.6 1.83 impa, real, adeq
meme 50 80.0 / 16.0 / 4.0 1.96 adeq, impa, real
ConservativeMemes 45 22.2 / 68.9 / 8.9 2.02 hoax, prop, real
Total 1,184 78.0 / 17.2 / 4.8 2.11 adeq, impa, hoax
Table 1: The number of memes (#m) in the top 6 frequent subreddits, along with percentages of convinced, skeptical, and neither stances, average number of involved frames (#f), and top 3 frequently used frames.
Filtering CC Memes.
Two master’s students in computational linguistics manually annotated all images to ensure a curated collection of CC memes: climate-associated and in the format of a meme. Annotators first assess the relevance of these images to climate change, retaining only samples where climate change was a central theme. They then identify whether a sample qualified as a meme by examining its combination of visual and textual elements, humorous or satirical intent, and relevance to cultural or social contexts. As Figure 2 Step 3 shows, tweets containing only text or lyrical statements paired with images are excluded.
3.2 Annotation
Stance Annotation
The SemEval 2016 shared task (Mohammad et al., 2016) introduced the stance detection task to classify tweets based on whether they are in favor of, against, or show neither stance towards specific targets, one of which was “Climate Change is a Real Concern.” We assess the stances of these 1,184 CC memes regarding whether the meme creators are convinced that climate change is real, skeptical, or neither (i.e., cannot tell), following terminologies from social science, particularly Hoffman (2011) (detailed in Appendix B.2).
Media Frame Annotation
In communication science, media frames are frequently identified to capture different, sometimes conflicting, perspectives on climate change. Jang and Hart (2015) propose five media frames to examine Twitter conversations on climate change. These frames include: , emphasizing whether the risk of climate change is present; hoax, questioning the faithfulness of public communication regarding the risk; cause, attributing the risk significantly to human activities; impact, highlighting the net negative consequences of the risk; and action, discussing necessary actions to address the risk. Ross and Rivers (2019) apply these five media frames to internet memes and exemplify the contrasting stances of individuals who are convinced of the CC issue and those who remain skeptical. Yet, they only present a handful of examples, and a dataset for quantitative analysis and modeling is still missing.
After adopting these five media frames and through six rounds of annotation revisions, we noticed the overly frequent occurrence of action. To provide a more fine-grained analysis of media frames on CC memes, we subdivide the action frame into the following four categories: allocation, propriety, adequacy, and prospect.
  • allocation captures discussions about the responsibility of certain groups, such as nations, organizations, or even generations, to take action on climate change than others;
  • propriety reflects debates on whether current actions are appropriate or effective;
  • adequacy highlights critiques regarding whether existing measures are sufficient to address climate risks or more actions are needed;
  • prospect explores perceptions of the potential outcomes of positive actions, distinguishing between climate doomists, who view catastrophe as inevitable, and climate risk realists, who believe meaningful prevention is still achievable Davidson and Kemp (2024).
This refinement also allows us to integrate additional dimensions observed in the data, most notably, moral evaluation (Entman, 1993; Gamson and Modigliani, 1989)—without introducing entirely separate frames. We embed moral reasoning within the allocation frame, which inherently concerns fairness, blame, and obligation, thereby capturing moral appeals in both skeptical and convinced memes while preserving theoretical clarity (see Appendix B.4 for guidelines and examples).
Inter-Annotator Agreement
The first author of this paper annotated stances and media frames on all 1,184 CC memes. To ensure the consistency of the annotations, we asked one master student in computational linguistics to annotate 200 randomly sampled memes following guidelines in Appendix B.2-B.3. We achieved high agreement for stance detection: 0.83 on Cohen’s Kappa. For media frame selection, since we allowed one or more labels per meme, we assess MASI distance and achieve an average score of 0.83. More, Cohen’s $\kappa$ for all eight frames exceeds 0.7 (see Appendix B.5).
3.3 OCR and Meme Caption
ClimateMemes includes two supplementary annotations: OCR correction and meme caption, as in Figure 2 Step 6. For each meme, we extract the embedded text via EasyOCR [^2] https://github.com/JaidedAI/EasyOCR and ask the two master students to correct any OCR errors manually. We follow Hwang and Shwartz (2023) and ask the annotators to write a concise caption describing the message that the meme conveys. We further investigate in §5 whether added explicit textual information helps stance and frame detection.
4 What Do Media Frames Reveal About Stance?
This section analyzes the interactions between stances and media frames in CC memes, including: ClimateMemes statistics (§4.1), frequently used media frames for convinced and skeptical memes (§4.2), concurrences of frames (§4.3), and whether specific frames signal a meme’s stance (§4.4).
4.1 ClimateMemes Statistics
Table 1 presents the number of memes in the top 6 frequent subreddits, along with their average number of frames and distribution of convinced, skeptical, and neither stances. About half of the 1,184 CC-associated memes are sourced from r/ClimateMemes, a community of climate activists. 94.1% memes from r/ClimateMemes exhibit a convinced stance, with the most frequently occurring frames being adequacy, cause, and impact. These frames discuss human activities as primary drivers of climate change, enumerate negative consequences, and call for more actions.
r/TheRightCantMeme, r/dankmemes each account for about 8% of the total memes, ranking second in tie. 83.5% of the memes from r/TheRightCantMeme demonstrate a skeptical stance, with the predominant frames being hoax, propriety, and cause. These frames reflect skepticism toward the truthfulness of the CC communications, the effectiveness of current actions, and the denial of human activity as the primary cause. In contrast, 82.3% of r/dankmemes memes exhibit a convinced stance, with being a common frame, highlighting that CC is indeed happening.
Despite continuous efforts to upsample skeptical memes and subreddits, ClimateMemes exhibits an unbalance where 78.0% memes are convinced and 17.2% are skeptical, most frequently employing adequacy and hoax frames, respectively (see Appendix C for a detailed frame distribution).
Figure 3: Monthly frequencies of media frames used in convinced versus skeptical memes.
4.2 Frame Preference
This subsection analyzes the framing preference of convinced versus skeptical stances over time. The publication time of 1,184 ClimateMemes memes spans eight years from March 2016 to September 2024. Figure 3 plots the monthly frequency of each frame separately for memes with convinced and skeptical stances from December 2018 to December 2024.333CC memes were quite rare before 2019. Two peaks occurred in September and December 2019, corresponding to Greta Thunberg’s speech at the United Nations Climate Summit and the COP25, for both convinced and skeptical memes. Interestingly, in convinced memes, the frequency of nearly all frames is significantly higher during these months, while in skeptical, only the hoax and propriety show an increase.
Figure 4 shows the probability of particular frames being involved in memes with convinced and skeptical stances. Among skeptical memes, 77.94% involve hoax, followed by propriety at 45.59%. Other frames appear in less than 15% memes. In contrast, frames in convinced memes are more diverse, with adequacy, impact, and cause being the most common, appearing in 42.1%, 40.20%, and 37.05% memes. Other frames, except for prospect, appear in 20%-30% memes.
Figure 4: Frame preference of convinced and skeptical memes.
4.3 Frame Concurrence
Since each meme can use multiple frames (2.11 frames/meme, cf. Table 1), Figure 5 investigates the concurrence of frames in convinced and skeptical memes. For skeptical, the concurrence of hoax and propriety frames is notably more potent than others. Rather, frame concurrences in convinced memes are more balanced across diverse combinations, similar to observations in Figure 4. Moreover, we notice that hoax has negative correlations with cause, impact, adequacy, and prospect, i.e., they tend not to co-exist (see Appendix D).
(a) convinced
(b) skeptical
Figure 5: Concurrence of media frames in convinced and skeptical memes.
4.4 Frame as a signal
Given that specific frames such as hoax are prevalent in skeptical memes, we examine whether frames serve as a good signal for stance detection. Figure 6 analyzes the likelihood of a meme being convinced or skeptical when a specific frame is used. We observe that when cause, impact, adequacy, and prospect appear in a meme, there is $>$80% probability that the meme holds a convinced stance. and allocation also appear more frequently in convinced memes. Conversely, hoax implies a 76.18% probability that the meme is skeptical, followed by propriety (59.87%).
To sum up, strategic media framing is essential in conveying stances in CC memes. Though hoax remains dominant in skeptical memes, framing is more diverse for convinced ones.
5 Stance and Media Frame Detection
To what degree can VLMs detect stance and frames in a meme, and how can we improve their performance? This section reports various experiments we performed on ClimateMemes.
5.1 Experimental Setups
Models.
We evaluate two open-source VLMs on multimodal memes: LLaVA-v1.6-Mistral-7B (LLaVA, Liu et al. 2024), Molmo-7B-D (Molmo,  Deitke et al. 2024), both of a visual encoder→cross-modal connector→LLM setup.
To compare, we experiment with text-only inputs on their LLM backbones: Mistral-7B (Mistral, Jiang et al. 2023) and Qwen2-7B (Qwen, Yang et al. 2024).
Figure 6: Stance distribution (in percentage) conditioned on frame usage in memes.
Data Partition.
We split ClimateMemes into train and test sets with an 8:2 ratio, and all models are evaluated on the 235 test memes.
Evaluation Scenarios.
In addition to zero-shot, we evaluate all models on $n$-shot experiments $n$ ranging from 1 to 4. Following Huang et al. (2024), we leverage relative sample augmentation to select top $n$ similar memes from train for each test meme based on the image and its human-corrected OCR. We also explore various input scenarios following Hwang and Shwartz (2023) to examine whether OCR, human, and synthetic meme caption can improve stance and media frame detection. We rotate stance and frame orders in prompts and report average over permutations (Zheng et al., 2023; Wang et al., 2024). For the backbone LLM baselines, we run experiments on text-only inputs.
Metrics.
We report accuracy and macro F1 for stance detection, focusing on the latter due to label imbalance. Since one or more media frames can be assigned to one meme, we binarily classify each frame and report the average over eight frames.
Model Inputs Stance Frame
Acc. F1 Acc. F1
baseline meme 80.85 29.80 73.83 43.98
LLaVA meme 77.31 39.08 51.87 45.63
meme+ocr 77.31 44.06 46.36 40.72
meme+syn 73.95 40.01 52.45 45.78
meme+syn+ocr 76.89 41.10 52.57 45.87
meme+hum 86.55 56.68 49.96 44.18
meme+hum+ocr 83.19 53.57 50.53 44.46
Molmo meme 47.06 28.16 60.37 52.60
meme+ocr 57.56 34.70 56.98 49.68
meme+syn 61.76 39.25 58.37 51.02
meme+syn+ocr 65.97 38.32 54.23 47.97
meme+hum 72.27 49.53 62.74 54.24
meme+hum+ocr 70.17 46.52 60.40 52.46
Mistral ocr 51.90 37.09 61.71 54.79
syn 58.23 36.06 59.03 53.01
syn+ocr 59.66 42.71 61.78 55.20
hum 79.32 60.54 64.61 58.31
hum+ocr 67.65 48.96 65.09 58.78
Qwen ocr 49.16 34.06 64.02 55.45
syn 68.91 44.66 60.33 53.98
syn+ocr 61.34 39.08 60.88 54.24
hum 73.11 53.28 65.86 58.23
hum+ocr 70.17 51.66 64.98 57.51
Table 2: Performance in accuracy and Macro-F1 on stance and frame detection with 4-shot setup. Backbone LLMs, Mistral and Qwen, only receive text input; syn = synthetic caption, hum = human caption. The baseline is calculated using majority vote, detail see Appendix E.
5.2 Which inputs help stance and frame detection in memes?
Table 2 examines how the number of shots and textual inputs influence VLM and LLM performances.
Zero-shot vs. Few-shot.
For both VLMs and their LLM backbones, few-shot setups outperform zero-shot on both tasks, evincing their in-context learning ability (0-4 shots in Appendix F).
VLMs vs. LLM backbones.
To what extent can visual inputs benefit VLM performances on meme understanding? While LLaVA has an edge over Mistral across various inputs on stance detection, both VLMs achieve lower scores on frame detection compared to LLMs. We hypothesize that VLMs are not pre-trained on meme datasets for frame detection. Yet, there already exists textual dataset related to framing Stede and Patz (2021); Frermann et al. (2023). It should also be noted that LLMs’ winning performances benefit from costly human annotations (OCR corrections [^4] We observed low-quality OCR; the average Levenshtein edit-distance before and after human correction is 60.75. and captions) or synthetic captions generated by VLMs. We also find that certain visual elements can mislead VLMs. Figure 7 shows two examples where LLaVA over-associates fire with . Though wildfires are often associated with CC existence, flames may also appear as metaphor or background.
(a) convinced stance with allocation and propriety frames
(b) skeptical stance with hoax frame
Figure 7: Two flame-related memes where LLaVA incorrectly predicts the Real frame.
OCR.
On stance detection, extra OCR input is beneficial for VLMs—though only in setups without human caption. For LLMs, feeding VLM-generated meme captions (syn) functions mostly better than using OCR, especially for Qwen. Combining OCR with synthetic captions can improve the scores for LLaVA in frame detection but always harms Molmo’s performance on both tasks. Importantly, OCR fails to help VLMs and LLMs further when combined with human captions. This underlines the importance of high-quality captions, leading to the overall best model for stance. Instead, for frames, LLMs outperform VLMs. We hypothesize that LLMs better grasp text inputs (especially captions) which aid fine-grained frame detection, while VLMs’ performance is lower on frames and benefits less from more explicit texts.
Human vs. Synthetic Caption.
Human meme captions improve performance on both tasks in almost all setups (except for frames with LLaVA). We leave it to future work to probe how meme captions help models understand stances and frames.
Frame #M LLaVA Molmo Mistral Qwen
Acc. F1 Acc. F1 Acc. F1 Acc. F1
44 30.69 26.84 68.99 59.43 46.60 43.88 72.13 60.32
hoax 81 51.60 46.90 71.54 59.43 71.33 60.54 71.01 61.46
caus 75 49.47 45.90 64.63 57.06 70.37 60.64 65.85 58.22
impa 77 48.03 45.37 60.59 54.64 71.33 61.29 61.54 56.03
allo 49 52.55 48.05 60.00 55.22 69.95 60.17 64.36 56.99
prop 80 53.62 49.54 56.86 53.18 64.10 56.65 64.10 57.57
adeq 81 56.86 50.50 59.68 54.84 64.26 56.98 65.21 58.28
pros 13 56.86 50.69 59.63 54.74 58.99 53.58 62.71 56.99
average 49.96 45.47 62.74 56.07 64.61 56.72 65.86 58.23
Table 3: Frame-specific performances with 4-shot meme+hum VLMs and hum LLMs. Best and worst scores per model are bolded and underlined. #M = number of test memes with the frame label.
5.3 Which frames are harder?
Table 3 reports per-frame performances of VLMs and LLMs. Consistent with overall performance, Molmo outperforms LLaVA in predicting all 8 frames. hoax achieves the highest score and propriety scores the lowest. For LLMs, Qwen outperforms Mistral with overall performance, but is not consistently the best model for all frames.
Model Base Input(+frame) $\Delta$Acc. $\Delta$F1
LLaVA meme -5.04 +0.45
meme+ocr -2.52 +5.26
meme+hum 0.00 +1.06
meme+hum+ocr -15.54 -7.03
meme+syn +2.52 +2.27
meme+syn+ocr -16.39 -4.74
Molmo meme +1.26 +2.81
meme+ocr +7.15 +3.33
meme+hum +5.46 +4.15
meme+hum+ocr +7.14 -0.30
meme+syn +6.31 +2.73
meme+syn+ocr +7.56 +3.11
Table 4: VLM performance changes on stance detection when gold frame labels are added as additional inputs.
5.4 Can frame labels help stance detection?
Table 4 investigates whether adding gold frame labels helps stance detection on 4-shot VLMs. Notably, for LLaVA (better at stance detection), incorporating frame information tends to reduce its performance. Instead, for Molmo (better at frame detection), adding frame information generally boosts its performance. Both models further improve their performance in the image and human caption setup. This suggests that stance and frame detection could benefit from multi-task training, improving performance through shared knowledge.
6 Meme Understanding through the Lens of Communication Science
To understand which aspects challenge models in meme understanding, we apply an interdisciplinary approach, integrating three concepts from communication science research: humor type, personalization, and responsibility. These concepts are critical in understanding the construction of meaning in affective climate communication on social media and may pose challenges for detection tasks. Humor is a key feature of memes as they are usually created with the intention to entertain people  Taecharungroj and Nueangjamnong (2015). Thus, we analyzed different humor types, such as puns, sarcasm, and surprise, which can have varying effects on readers. Personalization is a common communication strategy in political communication in general and that simplifies complex political issues by focusing on individual actors instead of groups.
We recruited two bachelor’s students in communication science to annotate humor type, person, and responsibility on 235 ClimateMemes test memes. Our guidelines are adapted from a comprehensive codebook on “Climate Change and Social Media” [^5] Anonymized version available at: https://osf.io/3hqdk?view_only=dd6035e7b03542e4a66c2fafa4bf0d7d allowing multiple labels per item (detailed definitions in Appendix G). Table 5 shows the most common labels, number of relevant memes, and LLaVA’s subsample performance.
In stance detection, LLaVA performs well on memes with humor types exaggeration and personification. Memes with irony and puns are difficult, receiving the lowest accuracy and F1. Memes with ordinary and celebrity under personalization categories show strong performance, but NGO stands out as challenging. For responsibility, memes concerning the macro-level are the easiest for the model, while meso-level memes are the hardest. Since micro/meso-level responsibilities address specific individuals (e.g., politicians) or organizations (e.g., companies), we hypothesize that macro-level responsibility (e.g., the society) leads to less variation and eases meme understanding.
In frame detection, memes with humor types pun and joke are the hardest. Under personalization, memes featuring NGO or celebrities are easier for the model than political ones. Frame detection on memes attributing responsibility mirrors results on stance detection, with macro performing best and meso worst. In sum, this interdisciplinary annotation using taxonomies from communication science provides insights into aspects that caused difficulties in meme stance and frame detection. Appendix H provides additional error analyses.
Concept Label #M Stance Frame
Acc. F1 Acc. F1
humor type irony 33 78.79 53.30 64.91 57.43
compare 25 88.00 69.90 65.75 54.67
surprise 21 85.71 50.42 64.58 56.82
personification 21 95.24 82.05 67.26 61.39
joke 19 84.21 52.53 58.14 50.86
exag 10 100.00 100.00 60.31 56.37
pun 5 80.00 33.33 59.69 42.86
Total 134 87.42 63.21 62.95 54.34
personalization ordinary 86 88.37 66.60 62.55 54.03
celebrity 25 88.00 53.23 61.75 52.27
politicial 14 85.71 78.79 54.80 47.93
NGO 14 50.00 37.18 66.74 55.43
Total 139 78.02 58.95 61.87 54.35
responsibility macro 50 88.00 51.06 64.88 58.39
meso 37 89.19 31.43 60.47 52.10
micro 37 83.78 46.88 60.26 52.56
Total 124 86.99 43.12 61.84 54.40
Table 5: Llava 4-shot meme+hum results on test subsamples with humor, person, and responsibility labels.
7 Conclusion
We introduce ClimateMemes, a new dataset of climate change memes annotated with stance and media frames. We demonstrate that media frame preferences are strong indicators of stance, with convinced and skeptical stances favoring distinct frames. We compare VLMs and LLMs and identify challenges in understanding multimodal information. Our interdisciplinary paper integrates concepts from communication science and reveals which meme aspects challenged the model.
Limitations
Potential Bias in the Sample Due to the Platform
Our dataset is exclusively composed of memes collected from Reddit, which introduces a potential bias. By focusing solely on this platform, we limit the diversity of content that could be found on other platforms like Twitter, Facebook, Instagram, or 4chan. Each platform has its own user base, culture, and way of sharing and discussing memes, which could result in differences in the types of memes that are shared. This platform-specific limitation means that our findings might not be fully representative of meme trends across the internet as a whole. The memes on Reddit, for instance, might reflect particular political, social, or cultural preferences that do not necessarily align with those on other platforms, making the sample less generalizable.
Distribution of Meme Stances and Frames Might Be Different on Other Platforms
The stance and framing of memes on Reddit may not reflect the patterns found on other social media platforms. Reddit has a unique structure, where specific subreddits cater to distinct interests, communities, and ideologies, which could influence the stances and frames adopted in memes. For example, some subreddits may have a higher concentration of memes that are either supportive or skeptical of climate change, while other platforms might exhibit different trends. Memes on Twitter or Instagram could carry different connotations, tones, or styles that might not be as prevalent on Reddit. Thus, the distribution of meme stances and frames could vary significantly across platforms, and a more comprehensive understanding of meme discourse would require analyzing multiple platforms to account for these differences.
Only a Single Annotator
We acknowledge the limitation of only a single annotator. Despite limited resources, we carefully refined our annotation guidelines through six iterations, totaling about 540 hours of annotation work.
Monthly Frequency: Sample Size May Be Too Small in Some Months to Derive Conclusions About Temporal Trends
The monthly frequency of memes in our dataset might not be large enough in certain months to allow for meaningful conclusions about trends or changes over time. If the sample size in a given month is too small, it becomes difficult to accurately detect shifts in meme stances, frames, or topics that may occur over longer periods. This limitation could obscure any subtle trends or variations in the frequency of specific meme types or themes, making it harder to assess how the discourse around a particular subject evolves. For instance, if a meme trend spikes during a specific event but the dataset contains very few memes from that month, it might not reflect the broader public sentiment or provide an accurate representation of the temporal dynamics.
Visual inputs for VLMs
We did not evaluate VLMs without visual input, and using the LLM backbone alone might not be 100% comparable to running a VLM without image input, because VLMs are fine-tuned on different datasets.
Ethics Statement
All annotations were conducted in accordance with ethical guidelines, ensuring that annotators were not exposed to any psychologically distressing content during the process. All annotators are paid according to national standards.
Acknowledgements
This work is supported by the KLIMA-MEMES project funded by the Bavarian Research Institute for Digital Transformation (bidt), an institute of the Bavarian Academy of Sciences and Humanities. The authors are responsible for the content of this publication.
Appendix A Subreddits in ClimateMemes
subreddit frequency description
ClimateMemes 591 The community to share environmental memes of prime quality. We advocate for climate action through funny captions and satire. Release your inner Greta, share your dankest decarbonization memes and raise global awareness to save the planet! Discuss climate strikes, climate change denial and doomerism, Fridays For Future, facts and news about nature, climate crisis quotes, ecology, Extinction Rebellion, and the end of the world.
TheRightCantMeme 91 Get your fix at left-wing Reddit alternatives: Hexbear and Lemmygrad. Also check out the Discord.
dankmemes 90 D A N K
memes 76 Memes! A way of describing cultural information being shared. An element of a culture or system of behavior that may be considered to be passed from one individual to another by nongenetic means, especially imitation.
meme 50 r/meme is a place to share memes. We’re fairly liberal but do have a few rules on what can and cannot be shared.
ConservativeMemes 45 Become a ConservativeMemes subscriber! — Click the JOIN button now, and post your Conservative Memes later at /r/ConservativeMemes !!! — If you like political humor, political memes, politically incorrect memes, or conservative memes, this is the sub for you!
PoliticalCompassMemes 39 Political Compass Memes
terriblefacebookmemes 30 Community for all those terrible memes your uncle posts on facebook
ConspiracyMemes 18 This subreddit is devoted to memes relating to all things conspiracy. Things are pretty laid back around here so all people are welcome. The moderators believe in free speech and try not to moderate comments or posts unless it is absolutely necessary.
Memes_Of_The_Dank 15 This is a meme subreddit. That should be obvious by now. Also, it is slowly recovering from bots, and that’s good.
libertarianmeme 14 For an end to democracy and tyranny. For more information about our ideology, check out the Mises Institute
HistoryMemes 11 A place for history memes about events over 20 years ago.
CommunismMemes 9 A place to share memes about communism.
PoliticalMemes 7 We’re striving for equality here. Not "equality" in the sense that we’ll allow people to post bigoted nonsense or perpetuate a false equivalency of entities, but "equality" in the sense that we are all co-inhabitants of this flying rock and need to learn to live together peacefully.
dank_meme 7 Dank Memes
MemeEconomy 7
sciencememes 7
PrequelMemes 6 Memes of the Star Wars Prequels.
PresidentialRaceMemes 6
MemeThatNews 6 Learn and comment on the news with memes.
AusMemes 6 The Australia Memes subreddit. Just waiting for a mate.
Animemes 6 A community for anime memes!
Marxism_Memes 5 MEMES ARE THE NEW PAMPHLETS JURY NULLIFICATION FOR COMRADE LUIGI!
TheLeftCantMeme 4 They make a lot of bad Political Memes
MinecraftMemes 3 A place to post memes about Minecraft! Our Discord Server can be found in the sidebar below.
AnarchyMemeCollective 3 A reddit for sharing anarchist memes and for discussing anarchism. If you share your own OC let us know and we may share if on our other platforms.
depression_memes 3 Memes about depression.
Funnymemes 3 "Where Laughter Lives: Your Daily Dose of the Funniest Memes!"
marvelmemes 2 Welcome to r/marvelmemes: The home of Marvel memes on Reddit!
lotrmemes 2 Come on in, have a seat! This subreddit is a warm resting place for all weary travelers who are fond of Tolkien and his works. We welcome all Tolkien related content! Grab a pint, a long pipe, and relax.
VegMeme 2 A place to share animal rights humor, cartoons, image macros etc, because if you can’t have a laugh at the hypocrisy and ignorance of carnists or have a good-natured laugh at ourselves you will probably become a misanthropic douchebag.
Jordan_Peterson_Memes 2 Welcome to the official subreddit for Jordan Peterson memes.
animememes 2 An anime meme subreddit that’s friendly for women, queer people, and generally marginalized anime fans who want a break from how toxic anime spaces usually are. Of course, anyone is welcome, but be respectful to the intention of the space.
AvatarMemes 2 A subreddit for memes and other humor related to the Avatar franchise. Jokes based on ATLA, LoK, etc. are welcome.
CoronavirusMemes 2 Opening back up due to popular demand, didn’t know people still wanted to post about the coronavirus. Monkeypoxmemes are allowed. Getting a laugh out of the Coronavirus while we still can, and spreading happiness in a time of distress.
SequelMemes 1 Memes of the Star Wars Sequels
VoluntaristMemes 1 Memes for voluntarists and other liberty loving people.
CommunistMemes 1 Communism is always the end goal!
SimpsonsMemes 1 Memes from The Simpsons!
MemePiece 1 The best place to find One Piece memes! We celebrate the comedic and casual side of the series One Piece. Casual or low effort content, normally removed from r/OnePiece, is likely welcome!
CrusadeMemes 1 DEUS VULT
MemeReserve 1 The Doomsday Global Meme Vault is a fail-safe meme storage sub, built to stand the test of time — and the challenge of natural or economical collapse. Only for the best memes!
GameOfThronesMemes 1 This subreddit is currently closed. Please check out r/aSongofMemesAndRage for memes based off GOT, ASOIAF, etc.
IncrediblesMemes 1 It’s showtime
memesITA 1 Pizza, pasta & memes.
AnimeMeme 1 AnimeMeme for anime memes.
YouBelongWithMemes 1 The official meme subreddit for r/TaylorSwift
Table 6: ClimateMemes’s 47 subreddits with their descriptions and meme frequency.
Table 6 details the public descriptions and meme frequencies of 47 subreddits in ClimateMemes.
Appendix B Annotation Guidelines
B.1 Filtering Climate Change Memes
Is this image associated with the topic of climate change?
Often images that discuss terms such as “climate change,” “global warming,” “greenhouse gas,” “carbon emission,” “fossil fuel,” “ozone,” “air pollution,” “carbon dioxide emissions,” “deforestation,” “industrial pollution,” “rising sea levels,” “extreme weather,” “melting glaciers,” “ocean acidification,” “biodiversity loss,” “ecosystem disruption,” “carbon capture,” “carbon storage,” “soil carbon,” “renewable energy,” “sustainable practices,” “Paris Agreement,” “Kyoto Protocol,” “carbon tax,” “emissions trading schemes,” “green technology,” “sustainable technology,” and “environmental change” are associated with climate change.
Additionally, if the meme features a well-known environmentalist or a political leader who has made statements related to climate change and environmental protection, it should also be considered as "associated with climate change." If you encounter an unfamiliar person, please use Google to search and confirm.
Is this image a meme? Is it a cartoon?
Memes are created by taking an existing widespread image and attaching new meaning to it by adding text within the image. A political cartoon, also known as an editorial cartoon, uses caricatures and satire to express an artist’s opinion on current events, often critiquing political leaders, social issues, or corruption through humor and exaggeration. A cartoon style often features exaggerated characters and simplified forms, and the text is usually in hand-drawn fonts that match the casual, expressive tone of the illustration. Both memes and political cartoons are considered memes in this study.
B.2 Stance Annotation
What is the stance of this CC meme?
We annotate the stances of CC memes into the following three categories: convinced, skeptical and neither.
  • convinced: Accepts environmental risks, supports regulation of harmful activities, and reflects egalitarian and communitarian values.
  • skeptical: Downplays or denies environmental risks, opposes regulation, and prioritizes individual freedom and commerce.
  • neither: Does not align with convinced or skeptical stance and may present a neutral or unrelated stance.
B.3 Media Frame Annotation
Climate change, a critical global issue, refers to long-term alterations in temperature and weather patterns, largely driven by human activities such as fossil fuel combustion. As this issue gains prominence, memes—images paired with text—have become a widespread tool for expressing opinions and social commentary online via media framing.
In this task, you will be given CC memes and will be asked the following question: which media frames are used in these CC memes? Choose one or multiple that apply.
  • emphasizes that there are evidences indicating that CC is occurring;
  • hoax questions the faithfulness of public communication by politicians, the media, environmentalists, etc., e.g., if they are misrepresented or manipulated;
  • cause attributes human activities as a significant cause of CC;
  • impact highlights that CC leads to more net negative outcomes than if there was no CC;
  • allocation captures discussions about the responsibility of certain groups, such as nations, organizations, or even generations, to take action on climate change than others;
  • propriety reflects debates on whether current actions are appropriate or effective;
  • adequacy highlights critiques regarding whether existing measures are sufficient to address climate risks or more actions are needed;
  • prospect explores perceptions of the potential outcomes of positive actions, distinguishing between climate doomists, who view catastrophe as inevitable, and climate risk realists, who believe meaningful prevention is still achievable Davidson and Kemp (2024).
B.4 Discussion about Moral Frame
While moral judgment is a salient feature in many memes—such as blaming past generations, exposing hypocrisy, or invoking responsibility for future generations—it is theoretically debated whether morality should be treated as a distinct media frame or as an underlying component of framing processes (Entman, 1993; Gamson and Modigliani, 1989). Rather than isolating morality as a standalone frame, we integrate moral reasoning into the allocation frame. This choice reflects our observation that moral claims are often embedded in discussions of responsibility and fairness, and enables us to capture moral stances consistently across both convinced and skeptical memes without losing theoretical coherence.
B.5 Frame-level IAA
Table 7 presents per-frame Cohen’s $\kappa$ for inter-annotator agreement (IAA) among two annotators.
Frame $\alpha$
Real 0.810
Hoax 0.868
Cause 0.825
Impact 0.711
Action_allocation 0.786
Action_propriety 0.777
Action_adequacy 0.740
Action_prospect 0.834
Table 7: Cohen’s $\kappa$ scores for IAA among two annotators.
Appendix C Frame Distribution in Climate Change Memes
Among the 1,184 climate change-associated memes we identified, we annotated a total of 2,499 frames, averaging approximately 2.11 frames per meme. Table 8 presents the distribution of these frames across eight main categories. The adequacy frame is the most frequently annotated, followed closely by hoax, impact, and causeframes. prospect is the least frequent frame, indicating fewer memes discussing future-oriented aspects of climate change.
Frame hoax cause impact allocation propriety adequacy prospect Total
Count 269 387 370 395 208 382 419 69 2,499
% 10.8% 15.5% 14.8% 15.8% 8.3% 15.3% 16.8% 2.8% 100%
Table 8: Distribution of frames annotated in the 1,184 climate change-associated memes (2,499 in total, averaging 2.11 per meme).
Appendix D Frame Correlation
We demonstrate the correlation among 8 frames in Figure 8.
Figure 8: Correlation heatmap of frames. The values represent pairwise Pearson correlation coefficients. Values marked with * indicate corresponding p-values less than 0.05, indicating significance.
Appendix E Stance and Frame Distribution in Test Set & Majority Vote Baseline
In our dataset, we analyzed a total of 235 cases annotated with stance labels. The distribution of the stances in these cases is summarized in Table 9.
Stance Convinced Skeptical Neither
Count 190 32 13
Table 9: Distribution of stance labels in the 235 test memes.
The majority class is Convinced with 190 cases. A majority-vote baseline predicting this class achieves an accuracy of 0.8085 and a macro F1 score of 0.2980.
Table 10 summarizes the distribution of “positive” and “negative” labels across the eight frames. For each meme, we computed a majority-vote baseline by predicts “negative” for all frames. The reported metrics are averaged over all 235 cases.
Frame Positive Negative Total
42 193 235
hoax 79 156 235
cause 75 160 235
impact 74 161 235
allocation 49 186 235
propriety 78 157 235
adequacy 82 153 235
prospect 12 222 235
Table 10: Distribution of “positive” and “negative” labels across eight frames in the 235 test memes.
This baseline, predicting “negative” for all frames, achieves an average accuracy of 0.7383 and an average macro F1 score of 0.4398 across the dataset.
Appendix F Additional Experimental Results
We show the full experimental results of stance detection of VLMs in Table 11, of LLMs in Table 12, frame detection of VLMs in Table 13, and of LLMs in Table 14. In line chart Figure 9, we also present VLM performances on stance and media frame detection with different shot and input setups.
Model #S Inputs Acc. F1 precision recall
LLaVA 0 meme 76.89 28.98 26.75 31.61
meme+ocr 45.80 30.80 35.87 37.73
meme+hum 68.91 46.12 58.29 55.05
meme+hum+ocr 66.39 44.83 48.50 55.75
meme+syn 55.46 34.41 36.67 38.22
meme+syn+ocr 56.72 35.75 38.32 40.48
1 meme 64.29 37.75 39.42 38.16
meme+ocr 60.08 42.27 41.78 49.04
meme+hum 73.95 55.68 54.30 63.42
meme+hum+ocr 71.85 52.60 53.35 61.04
meme+syn 62.18 37.19 37.69 40.12
meme+syn+ocr 66.39 41.85 41.85 44.24
2 meme 68.07 34.77 40.36 35.37
meme+ocr 68.91 43.26 42.31 45.27
meme+hum 82.35 56.08 55.79 58.62
meme+hum+ocr 74.37 45.31 43.33 52.29
meme+syn 70.59 41.25 41.07 42.70
meme+syn+ocr 73.95 43.79 44.55 44.95
3 meme 73.95 35.64 45.55 35.40
meme+ocr 74.79 45.90 45.97 45.95
meme+hum 83.61 54.21 54.73 55.01
meme+hum+ocr 81.09 51.65 52.37 53.97
meme+syn 73.11 40.04 38.67 42.22
meme+syn+ocr 76.47 41.10 39.71 42.73
4 meme 77.31 39.08 49.55 40.69
meme+ocr 77.31 44.06 46.19 42.86
meme+hum 86.55 56.68 62.50 54.48
meme+hum+ocr 83.19 53.57 55.13 53.97
meme+syn 73.95 40.01 38.76 41.69
meme+syn+ocr 76.89 41.10 39.60 42.90
Molmo 0 meme 42.02 25.02 30.30 30.31
meme+ocr 39.08 24.32 31.31 31.70
meme+hum 47.06 33.27 40.09 46.07
meme+hum+ocr 43.70 31.50 40.16 45.55
meme+syn 30.67 25.91 37.67 41.08
meme+syn+ocr 28.99 24.80 37.18 40.38
1 meme 61.76 30.83 30.98 32.34
meme+ocr 52.10 31.68 34.73 37.93
meme+hum 59.66 41.76 43.43 51.90
meme+hum+ocr 53.78 37.00 41.04 48.83
meme+syn 53.78 37.45 39.43 46.01
meme+syn+ocr 56.72 39.45 41.55 48.08
2 meme 45.38 26.90 31.27 32.56
meme+ocr 52.52 31.22 34.30 36.36
meme+hum 68.49 47.99 46.59 55.31
meme+hum+ocr 66.39 43.66 44.16 52.27
meme+syn 57.38 35.00 36.62 38.26
meme+syn+ocr 63.03 36.54 36.65 41.55
3 meme 47.06 28.81 33.69 37.59
meme+ocr 55.88 33.20 35.32 39.48
meme+hum 68.49 47.99 47.64 53.57
meme+hum+ocr 68.49 44.41 45.90 51.40
meme+syn 60.34 37.13 37.96 41.21
meme+syn+ocr 63.03 35.93 36.24 40.68
4 meme 47.06 28.16 32.95 34.98
meme+ocr 57.56 34.70 36.37 42.78
meme+hum 72.27 49.53 50.21 55.34
meme+hum+ocr 70.17 46.52 52.67 54.70
meme+syn 61.76 39.25 40.45 45.16
meme+syn+ocr 65.97 38.32 37.73 43.63
Table 11: VLMs’ performance in terms of accuracy and Macro-F1 on stance detection. Hum = human caption, syn = synthetic caption, #S = number of shots.
Model #S Inputs Acc. F1 precision recall
Mistral 0 ocr 46.64 33.90 37.53 42.64
hum 66.67 51.62 50.70 65.00
hum+ocr 62.71 47.90 47.81 60.98
syn 39.41 29.94 37.35 40.45
syn+ocr 41.53 30.61 36.59 41.32
1 ocr 48.10 35.28 39.01 44.88
hum 69.20 50.68 49.24 59.52
hum+ocr 62.87 49.07 48.56 64.09
syn 54.62 37.70 39.04 44.40
syn+ocr 52.94 40.13 41.31 51.75
2 ocr 48.74 36.19 39.70 45.89
hum 71.73 52.68 50.88 60.56
hum+ocr 63.45 46.47 46.22 56.28
syn 56.78 39.92 40.15 46.82
syn+ocr 53.81 42.10 43.23 56.23
3 ocr 52.32 38.59 40.81 49.21
hum 77.54 57.36 55.22 62.96
hum+ocr 69.33 50.81 49.02 59.57
syn 55.93 36.06 37.34 41.03
syn+ocr 57.98 43.90 43.68 55.56
4 ocr 51.90 37.09 39.52 46.43
hum 79.32 60.54 57.57 66.95
hum+ocr 67.65 48.96 47.40 57.36
syn 58.23 36.06 37.24 40.24
syn+ocr 59.66 42.71 42.97 51.47
Qwen2 0 ocr 41.95 31.49 36.80 45.72
hum 64.29 50.29 55.11 59.46
hum+ocr 67.51 49.31 49.53 59.05
syn 55.88 36.20 53.12 51.88
syn+ocr 61.76 37.06 43.43 47.13
1 ocr 43.28 31.34 35.86 40.83
hum 64.71 48.60 50.08 58.11
hum+ocr 65.13 49.02 50.31 59.80
syn 61.34 40.18 42.22 43.47
syn+ocr 65.13 39.49 42.33 39.37
2 ocr 44.96 32.74 37.72 44.56
hum 70.59 52.60 55.23 59.66
hum+ocr 69.33 50.69 52.38 61.32
syn 62.61 40.74 50.80 43.77
syn+ocr 64.29 42.51 47.10 46.20
3 ocr 47.06 34.10 38.19 42.38
hum 73.53 55.19 57.97 61.74
hum+ocr 66.81 49.35 50.24 58.10
syn 62.18 40.97 47.24 44.47
syn+ocr 59.24 39.80 44.51 45.65
4 ocr 49.16 34.06 37.96 39.98
hum 73.11 53.28 55.79 58.30
hum+ocr 70.17 51.66 52.22 61.01
syn 68.91 44.66 47.97 49.62
syn+ocr 61.34 39.08 40.29 44.12
Table 12: LLMs’ performance in terms of accuracy and Macro-F1 on stance detection. Hum = hum caption, syn = synthetic caption, #S = number of shots.
Model #S Inputs Acc. F1 precision recall
LLaVA 0 meme 46.27 41.06 44.59 50.60
meme+ocr 44.65 40.27 46.50 50.39
meme+hum 49.38 45.18 51.77 53.19
meme+hum+ocr 46.70 43.10 51.23 52.08
meme+syn 44.93 41.34 50.59 50.22
meme+syn+ocr 44.19 40.85 50.65 50.35
1 meme 55.86 47.82 50.96 53.05
meme+ocr 53.09 45.94 49.77 52.87
meme+hum 56.42 49.85 53.13 56.16
meme+hum+ocr 56.62 49.44 53.21 55.01
meme+syn 57.71 49.35 52.78 53.97
meme+syn+ocr 56.97 48.89 52.42 53.85
2 meme 54.87 47.03 50.80 52.67
meme+ocr 50.95 44.22 47.79 52.40
meme+hum 55.86 48.91 52.19 55.72
meme+hum+ocr 54.85 47.80 51.38 54.34
meme+syn 58.13 49.56 52.88 53.88
meme+syn+ocr 57.13 48.74 52.03 53.51
3 meme 53.44 46.15 49.55 53.47
meme+ocr 48.86 42.50 45.73 53.41
meme+hum 51.91 45.14 47.76 55.38
meme+hum+ocr 53.56 46.75 50.18 55.09
meme+syn 55.62 47.85 51.38 54.45
meme+syn+ocr 55.53 47.74 51.29 54.54
4 meme 51.87 45.63 50.22 53.83
meme+ocr 46.36 40.72 43.54 53.40
meme+hum 49.96 44.18 47.60 55.55
meme+hum+ocr 50.53 44.46 48.10 54.18
meme+syn 52.45 45.78 49.53 54.70
meme+syn+ocr 52.57 45.87 49.43 54.81
Molmo 0 meme 49.38 43.89 48.00 55.04
meme+ocr 53.90 46.13 49.03 55.52
meme+hum 53.66 47.71 51.30 56.43
meme+hum+ocr 53.21 47.23 50.97 56.12
meme+syn 51.80 45.80 49.40 55.56
meme+syn+ocr 52.77 46.20 49.21 55.35
1 meme 54.88 47.73 50.06 55.65
meme+ocr 58.20 48.36 50.89 55.82
meme+hum 57.75 50.68 54.44 57.07
meme+hum+ocr 58.32 50.04 53.30 57.14
meme+syn 50.51 44.61 49.13 54.24
meme+syn+ocr 51.42 45.30 49.80 55.25
2 meme 53.58 46.62 48.77 54.85
meme+ocr 54.89 47.35 51.36 55.12
meme+hum 59.84 52.15 55.45 57.88
meme+hum+ocr 56.88 49.58 53.43 56.63
meme+syn 52.73 46.66 51.48 55.47
meme+syn+ocr 54.26 47.58 51.83 55.43
3 meme 55.72 48.61 51.00 55.14
meme+ocr 54.02 47.28 51.17 56.01
meme+hum 60.04 52.05 54.70 57.93
meme+hum+ocr 58.36 50.93 54.35 57.98
meme+syn 56.16 49.14 53.04 55.84
meme+syn+ocr 53.71 47.00 50.81 55.44
4 meme 60.37 52.60 54.99 57.41
meme+ocr 56.98 49.68 53.12 56.63
meme+hum 62.74 54.24 56.65 59.35
meme+hum+ocr 60.40 52.46 55.06 58.54
meme+syn 58.37 51.02 54.12 57.04
meme+syn+ocr 54.23 47.97 51.53 56.92
Table 13: VLMs’ performance in terms of accuracy and Macro-F1 on frame detection. Hum = human caption, syn = synthetic caption, #S = number of shots.
Model #S Inputs Acc. F1 precision recall
Mistral 0 ocr 55.03 49.79 54.96 55.65
hum 58.72 53.35 57.91 58.97
hum+ocr 57.73 52.46 57.30 58.47
syn 53.52 48.26 53.95 55.06
syn+ocr 54.70 49.30 54.48 55.68
1 ocr 56.32 50.75 55.43 56.04
hum 60.81 55.08 59.07 59.85
hum+ocr 59.79 54.18 58.18 59.29
syn 54.97 49.76 54.58 55.17
syn+ocr 56.98 51.78 56.35 57.14
2 ocr 59.18 52.51 55.95 56.65
hum 62.38 56.23 59.36 60.41
hum+ocr 62.27 55.96 58.96 60.03
syn 56.44 50.65 55.15 55.28
syn+ocr 58.61 52.43 56.00 56.54
3 ocr 60.55 53.80 57.04 57.88
hum 63.67 57.55 60.39 61.55
hum+ocr 63.84 57.63 60.16 61.78
syn 58.45 52.48 56.41 56.81
syn+ocr 60.45 54.12 57.55 58.25
4 ocr 61.71 54.79 57.59 58.68
hum 64.61 58.31 61.00 62.31
hum+ocr 65.09 58.78 61.06 62.94
syn 59.03 53.01 56.42 57.48
syn+ocr 61.78 55.20 58.10 59.12
Qwen2 0 ocr 56.60 49.47 53.29 54.49
hum 65.07 55.32 57.85 58.63
hum+ocr 60.82 52.94 55.81 56.99
syn 54.24 47.82 52.28 53.33
syn+ocr 53.69 47.41 51.66 53.06
1 ocr 59.59 51.82 54.94 55.63
hum 64.07 56.54 59.53 60.69
hum+ocr 63.39 55.68 58.50 59.15
syn 57.55 50.92 54.79 54.84
syn+ocr 57.69 51.08 54.90 55.30
2 ocr 62.11 54.01 56.95 57.72
hum 64.84 57.34 59.96 61.24
hum+ocr 64.31 56.57 58.88 60.05
syn 58.36 52.09 55.99 57.10
syn+ocr 60.65 53.87 57.07 58.42
3 ocr 63.14 54.32 56.92 57.65
hum 65.76 58.22 60.59 61.94
hum+ocr 65.32 57.52 59.76 61.05
syn 59.85 53.24 56.58 57.65
syn+ocr 60.85 54.16 57.57 58.96
4 ocr 64.02 55.45 57.99 58.80
hum 65.86 58.23 60.66 62.18
hum+ocr 64.98 57.51 60.01 61.12
syn 60.33 53.98 57.27 58.26
syn+ocr 60.88 54.24 57.45 58.45
Table 14: LLMs’ performance in terms of accuracy and Macro-F1 on frame detection. Hum = human caption, syn = synthetic caption, #S = number of shots.
(a) Stance detection
(b) Frame detection
Figure 9: Accuracy and Macro-F1 of VLMs on stance and media frame detection with different shot and input setups.
Appendix G Definition of CC-associated communication science concepts
Humor Type.
For the humor types category, the content format used to create humor in memes is coded. Humor types are initially independent of the respective humor style. Following (Taecharungroj and Nueangjamnong, 2015), a distinction is made between seven humor types, several of which can in principle be used simultaneously in a meme.
  • Puns use language to construct new meanings or use words or phrases in a way that suggests two interpreta-tions, e.g. words that are pronounced the same but have different meanings.
  • Personifications (personification) are used when human s and/or behavior are attributed to other objects such as animals, plants or ob-jects.
  • Exaggerations and understatements are disproportionate enlargements or reductions of a fact or context. Something is depicted as being larger or smaller than it (supposedly) actually is. Both the behavior of people and the consequences of events are depicted larger or smaller.
  • Comparisonsare combinations of two or more elements (e.g. before and after pictures) to construct a funny situa-tion.
  • Irony and sarcasm refers to the use of words to express the opposite of what one actually means.
  • Surprise is the use of unexpected elements in memes. Memes with this element have a surprising ending/resolution.
  • Jokes and nonsense describes content with no particular meaning and non-serious statements or actions that are only in-tended to make us laugh.
Personalization.
Who is shown in the picture?
  • Political actors include heads of state, members of government, official state delegates to the COP, ministers, representatives of institutions such as the UN or EU.
  • NGO members or environmental activists. Members of protest movements such as Fridays for Future are considered environ-mental activists, whereas “normal” participants in demonstrations are coded as “normal citizens”.
  • Celebrities are famous people who do not have an official political function. This includes, for example, people such as athletes, actors, influencers or artists.
  • Normal or ordinary citizens are people who are not clearly assigned to one of the other categories.
Responsibility.
To whom the responsibility for solving or combating the climate problem is attributed. The aim is to record who should take measures against climate change (e.g. more environmental protection, fewer emissions) or who isexpected to do so. Responsibility can be explicitly attributed or suggested by listing necessary measures that only a certain group can take.
  • Responsibility at micro level: Responsibility for individual persons such as politicians, activists, entrepreneurs, etc.
  • Responsibility at meso level: Responsibility for individual companies, institutions, parties, parliaments, governments.
  • Responsibility at macro level: Responsibility for certain countries, politics in general, the economic system, society, us as humanity, etc
Appendix H Case Study: Memes with Features in Communication
We present a case study on selected memes to analyze the model’s stance and frame prediction performance. Table 15 showcases three representative examples where our model made prediction errors. For each meme, we provide the associated features, gold and predicted stances and frames, as well as a detailed interpretation generated by the model. This qualitative analysis helps illustrate common challenges and nuanced aspects of meme communication that contribute to model misclassification.
Meme 1 Meme 2 Meme 3
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Features responsib_micro, humortypes_pun responsib_macro humortypes_irony
Gold Stance Convinced Skeptical Convinced
Predicted Stance Neither Convinced Skeptical
Gold Frames Allocation, Prospect Adequacy Hoax, Propriety, Adequacy
Predicted Frames Real, Impact Real, Hoax, Adequacy Real, Impact
Model Interpretation This meme humorously highlights the common experience of getting unexpected looks from strangers during St. Patrick’s Day celebrations. The image cleverly combines two related scenarios: wearing green for the holiday and taking public transportation. By showing muscular arms representing these two situations joining together, it suggests that both activities share a common outcome - drawing attention or being stared at. The meme taps into the relatable discomfort of standing out in a crowd, particularly during a holiday known for its vibrant green attire and large gatherings. Its simplicity and relatability make it an effective piece of internet humor that resonates with many people’s St. Patrick’s Day experiences. The meme features a man holding a dollar bill, with a caption that reads, "Is there a tax I can pay to stop COVID-19… Or does that only work for climate change?" This meme is a humorous take on the perceived difference in public opinion and government response to the COVID-19 pandemic and climate change. The implication is that people are more willing to pay taxes to address climate change, but not to combat the COVID-19 pandemic. The meme plays on the idea that some individuals prioritize environmental issues over public health crises, and it uses the man’s expression to convey a sense of frustration or disbelief at this perceived attitude. The humor arises from the juxtaposition of the man’s action (offering to pay a tax) with the implication that such an action would be more effective for climate change than for COVID-19. The meme features a series of images of a woman with a surprised or shocked expression, holding a bottle of ”ZERO“ with the text ”Pimples? Zero“ and ”Blackheads? Zero“ above each image. The humor in the meme comes from the exaggerated and unexpected response to the common skin issues of pimples and blackheads. The woman’s expression and the repeated use of the word ”zero“ suggest that the product being advertised is highly effective in eliminating these skin problems. The meme plays on the idea that the product is so powerful that it can completely eradicate these issues, which is an exaggerated and humorous claim. The meme is likely intended to be a parody of skincare product advertising, using hyperbole and visual cues to create a comedic effect.
Table 15: Qualitative examples of meme stance and frame prediction with model errors and interpretation generated by LLaVA.
Appendix I Elaboration About Use of AI Assistants
The authors used ChatGPT to polish writings for clarity and coherence and to assist with code generation. The authors manually inspected all ChatGPT suggestions and made corrections when necessary.